# Multiparty equality in the local broadcast model

### v1

@@ 1. Introduction
We consider the multiparty equality problem in graphs, where every vertex of a graph $G$ is given a $k$-bit input, and the goal of the vertices is to decide whether all inputs are the same. Each vertex can communicate with its neighbors in the graph, and the cost of the protocol is the total number of bits transmitted during the communication phase. After this phase, each vertex either accepts or rejects the instance; if every vertex was assigned the same input, then all vertices must accept the instance, and if two of the inputs differ then at least one vertex must reject the instance. The complexity of the multiparty equality problem is the minimum cost of a protocol solving the problem. Note that we only consider protocols that are deterministic and static, in the sense that the set of vertices sending messages and the size of their messages is independent of their input strings (they only depend on $G$ and $k$). One the other hand, the content of the messages is allowed to depend on the input strings.
@@
In the classical point-to-point communication model, each vertex communicates with its neighbors on different channels (one channel per vertex), so that sending the same message of $\ell$ bits to $d$ neighbors costs $d\ell$ bits of communication. A natural lower bound on the complexity of the multiparty equality problem in this model can be obtained by observing that at least $k$ bits need to be sent through every cut of the graph in every protocol. As a consequence, the complexity is at least $k$ times the minimum fractional transversal of cuts in the graph (or by duality, the maximum fractional packing of cuts), see [CR15, AES17]. Alon, Efremenko and Sudakov [AES17] showed that this linear programming lower bound can be attained asymptotically for a number of graph classes, including Hamiltonian graphs, and moreover they gave a protocol of cost within a factor $4/3$ of the optimal for every graph.
@@
Much less is known in the local broadcast communication model, which is the main focus of this paper. In this model, each message sent by a vertex $v$ is transmitted to all the neighbors of $v$, and the number of bits of this message is counted only once in the complexity of the protocol. The total cost of a protocol $\Pi$ on $G$, denoted $\mathsf{cost}_{\Pi}(G,k)$, is the sum of the number of bits broadcast by each vertex. The total complexity (minimum cost of a protocol solving the problem) of the multiparty equality problem in the local broadcast communication model on $G$ with $k$ bit inputs is denoted by $\operatorname{\mathsf{OPT}}(G,k)$. We will be mostly interested in the per-bit complexity $\operatorname{\mathsf{OPT}}(G)=\lim_{k\to\infty}\tfrac{1}{k}\operatorname{\mathsf{OPT}}(G,k)$, whose existence directly follows from Fekete’s subadditive lemma. Similarly, we define the per-bit cost of a protocol $\Pi$ as $\mathsf{cost}_{\Pi}(G)=\lim_{k\to\infty}\tfrac{1}{k}\mathsf{cost}_{\Pi}(G,k)$ (this limit does not exist for all protocols, but it does in all the protocols we consider in the paper). As in the point-to-point communication model, a simple linear programming lower bound on the complexity of the problem can be obtained by noting that for every cut, the sum of the number of bits broadcast by the vertices incident to the cut has to be at least $k$. This implies that a natural lower bound on $\operatorname{\mathsf{OPT}}(G)$ is the maximum fractional packing (or minimum fractional transversal) of boundaries in the graph $G$, where a boundary is the set of vertices incident to the edges of a given cut in the graph.
@@
We then design a new protocol in 2-connected graphs, which outperforms Theorem 1.2 on a number of natural graph classes. Our protocol uses ideas introduced by Alon, Efremenko and Sudakov [AES17] in the point-to-point model, relying on classical constructions in extremal combinatorics. The efficiency of our protocol depends on the minimum size of a total vertex cover in the graph $G$, denoted by $\mathsf{tvc}(G)$, where a total vertex cover is a subset of vertices which is both a vertex cover and a total dominating set. We prove the following result.
@@
Theorem 1.3 motivates the study of total vertex covers in graphs, for which we obtain new results which might be of independent interest. Our first application concerns the class of cycles. The linear programming lower bound on $\operatorname{\mathsf{OPT}}(C_{n})$ is of order $n/3$ (assigning weight $\tfrac{1}{3}$ to each vertex). On the other hand, as $\mathsf{wds}(C_{n})=\lfloor n/2\rfloor$, the simple protocol of Theorem 1.1 has per-bit cost $\lfloor n/2\rfloor$, and it can be checked that the protocol of Theorem 1.2 has per-bit cost $\tfrac{2}{3}n-\tfrac{2}{3}$ or $\tfrac{2}{3}n-\tfrac{1}{3}$, depending on the parity of $n$. As an immediate application of Theorem 1.3, we obtain a quasi-optimal protocol for cycles.
@@
It was proved in [CG17] that every cubic graph on $n$ vertices has a total vertex cover on a most $3n/4$ vertices. We extend this result to every degree by proving that for $d\geq 3$, every $d$-regular graph on $n$ vertices has a total vertex cover on a most $\tfrac{d}{d+1}\cdot n$ vertices. This is optimal for every $d$, since any total vertex cover of the complete graph $K_{d+1}$ (which is $d$-regular) contains at least $d$ vertices. It was proved in [CG17] that there are two more extremal connected graphs in the case $d=3$, and all the other connected cubic graphs have a total vertex cover with less than $3n/4$ vertices. Our result below implies that for $d\geq 5$, complete graphs are the only extremal examples.
@@ 2.1. Notation
In this paper, $\log$ denotes the binary logarithm. Let $G$ be a graph. For two disjoint sets $S,S^{\prime}\subseteq V(G)$, we denote by $E(S,S^{\prime})$ the set of edges of $G$ with the one endpoint in $S$ and the other one in $S^{\prime}$, and the cut defined by $S$ is $E(S,\bar{S})$, where $\bar{S}=V(G)\setminus S$. The boundary $B(S)$ of the cut defined by $S$ consists of all vertices incident to $E(S,\bar{S})$. Observe that for every $v\in V(G)$, the trivial boundary $B(\{v\})$ is precisely the closed neighborhood $N[v]$ of $v$ in $G$.
In the remainder we only consider connected graphs (otherwise the problem we consider has no solution, as vertices in different components cannot communicate to verify that their inputs are equal).
$2.$ A linear programming lower bound
@@ 4.1. The new protocol
Recall that a total vertex cover in a graph $G$ is a vertex subset that is both a total dominating set and a vertex cover. The minimum size of such a set is denoted by $\mathsf{tvc}(G)$. In this section we prove Theorem 1.3, restated below as Theorem 4.1.
@@
In terms of transversals, a total vertex cover is an (integral) transversal of open neighborhoods (total dominating set) and of edges (vertex cover). As we are dealing with graphs with no isolated vertices, a total vertex cover always exist, for instance every vertex cover of such a graph $G$ is dominating and can be made total by adding a neighbor of each of its elements, so $\mathsf{tvc}(G)\leq 2\mathsf{vc}(G)$.
@@
Let $H$ be a 2-connected $n$-vertex graph and let $m$ be a positive integer. Then there is a faithful host for $H$ with classes of vertices $U_{1},\dots,U_{n}$, each of size nmnm, containing a least $m^{2}/e^{10\sqrt{\log m\log n}}$ special copies of $H$.
@@
Let $n$ denote the number of vertices of $G$, that we call $v_{1},\dots,v_{n}$, and let $m$ be the minimum integer such that $\frac{m^{2}}{e^{10\sqrt{\log m\log n}}}\geq 2^{k}$. By Lemma 4.3 there is a faithful host $F$ for $G$ with classes of vertices $U_{1},\dots,U_{n}$, each of size nmnm, containing a least $2^{k}$ special copies of $G$. For every $i\in\left\{1,\dots,n\right\}$, let us name $u_{i}^{1},\dots,u_{i}^{nm}$ the vertices of $U_{i}$. We associate to each $k$-bit word $w$ a special copy $G_{w}$ of $G$ in $F$ in an injective way (i.e. if $w\neq w^{\prime}$ then $G_{w}\neq G_{w^{\prime}}$). Since there are at least $2^{k}$ special copies, this is always possible.
@@
Communication: Each vertex $v$ of $S$ broadcasts its identity to its neighbors and the other vertices remain silent.
Decision: A vertex $v$ of $G$ accepts the instance if and only if all identities received from its neighbors are consistent with its own identity.
@@ 4.2.3. Hypercubes
components, as desired. $\blacksquare$
@@ 4.2.4. Grids
In the $n\times n$ square grid $G=G_{n}$, consider the set $S$ of vertices $(i,j)\in[n]^{2}$ which either lie on the outerface, or are such that $i\in\{3j-2,3j-1\}\pmod{5}$ (in words, in row $j$ of the grid, we add to $S$ all vertices located in columns whose index is 3j23j-2 or 3j13j-1 modulo 5). Let $G_{S}$ be the subgraph of $G$ induced by the edges incident to $S$ (see Figure 1 for an illustration). Observe that $G_{S}$ is a spanning subgraph of $G$ and is 2-connected. By definition of $G_{S}$, it has a $S$ as total vertex cover so $\mathsf{tvc}(G_{S})\leq|S|\leq\tfrac{2}{5}\,n^{2}+4n$. By Corollary 4.2, the protocol $\Pi_{2}$ has per-bit cost at most $\tfrac{1}{2}\,|S|\leq\tfrac{1}{5}\,n^{2}+2n$. By Lemma 2.2 we have $\operatorname{\mathsf{OPT}}(G)\geq\tfrac{1}{5}\,n^{2}$, and the result follows. ∎
@@ 4.2.5. Regular graphs
Every cubic graph on $n$ vertices has a total vertex cover on a most $3n/4$ vertices.
@@
By extending the proof of Theorem 4.9 we can prove that for $d\geq 3$, every $d$-regular graph on $n$ vertices has a total vertex cover on a most $\tfrac{d}{d+1}\cdot n$ vertices. This is optimal for every $d$, since any total vertex cover of the complete graph $K_{d+1}$ (which is $d$-regular) contains at least $d$ vertices. It was proved in [CG17] that there are two more extremal connected graphs in the case $d=3$, and all the other connected cubic graphs have a total vertex cover with less than $3n/4$ vertices. The result below implies that for $d\geq 5$, complete graphs are the only extremal examples.
@@ 5.2. Open problems
A natural problem is whether the fractional lower bound $\tau_{\rm{bnd}}^{*}(G)$ can always be attained by $\operatorname{\mathsf{OPT}}(G)$. We have not been able to find a graph for which the two parameters provably differ. A natural class to consider for potential counterexamples is the class of 2-edge-connected cubic graphs, which has a clean expression for $\tau_{\rm{bnd}}^{*}$. For this class we have only been able to find protocols of per-bit cost $3n/8$, while the fractional lower bound is $n/4$.
@@
We have see that for various types of grids we could obtain asymptotically optimal protocols for equality, using total vertex covers in some 2-connected subgraphs, of size at most twice the fractional lower bound. The three types of grids we have considered can all be seen as Cayley graphs of $\mathbb{Z}^{2}$, with different generating sets. As the three constructions are very similar, a natural question is whether our constructions can be generalized to any Cayley graph of $\mathbb{Z}^{2}$.
Finally, it remains an open question whether the upper bound of $4\operatorname{\mathsf{OPT}}(G)$ on the cost of the protocol $\Pi_{1}$ in Theorem 1.2 can be improved.

### v2

@@ 1. Introduction
We consider the multiparty equality problem in graphs, where every vertex of a graph $G$ is given a $k$-bit input, and the goal of the vertices is to decide whether all inputs are the same. Each vertex can communicate with its neighbors in the graph, and the cost of the protocol is the total number of bits transmitted during the communication phase. After this phase, each vertex either accepts or rejects the instance; if every vertex was assigned the same input, then all vertices must accept the instance, and if two of the inputs differ then at least one vertex must reject the instance. The complexity of the multiparty equality problem is the minimum cost of a protocol solving the problem. Note that we only consider protocols that are deterministic and static [AES17, KV21, LV11], in the sense that the set of vertices sending messages and the size of their messages is independent of their input strings (they only depend on $G$ and $k$). One the other hand, the content of the messages is allowed to depend on the input strings. See [AES17] for a discussion on non-static protocols.
@@
In the classical point-to-point communication model, each vertex communicates with its neighbors on different channels (one channel per vertex), so that sending the same message of $\ell$ bits to $d$ neighbors costs $d\ell$ bits of communication. A natural lower bound on the complexity of the multiparty equality problem in this model can be obtained by observing that at least $k$ bits need to be sent through every cut of the graph in every protocol. As a consequence, as proved in [CR15, AES17], the complexity is at least $k$ times the minimum fractional transversal of cuts in the graph (or by duality, the maximum fractional packing of cuts), see Section 2.2 for more details. Alon, Efremenko and Sudakov [AES17] showed that this linear programming lower bound can be attained asymptotically for a number of graph classes, including Hamiltonian graphs, and moreover they gave a protocol of cost within a factor $4/3$ of the optimal for every graph. To the best of our knowledge, it is unknown whether computing the optimal cost of a protocol can be done in polynomial time.
@@
Much less is known in the local broadcast communication model, which is the main focus of this paper. In this model, each message sent by a vertex $v$ is transmitted to all the neighbors of $v$, and the number of bits of this message is counted only once in the complexity of the protocol. The total cost of a protocol $\Pi$ on $G$, denoted $\mathsf{cost}_{\Pi}(G,k)$, is the sum of the number of bits broadcast by each vertex. The total complexity (minimum cost of a protocol solving the problem) of the multiparty equality problem in the local broadcast communication model on $G$ with $k$ bit inputs is denoted by $\operatorname{\mathsf{OPT}}(G,k)$. We will be mostly interested in the per-bit complexity $\operatorname{\mathsf{OPT}}(G)=\lim_{k\to\infty}\tfrac{1}{k}\operatorname{\mathsf{OPT}}(G,k)$, whose existence directly follows from Fekete’s subadditive lemma [WIK26]. Similarly, we define the per-bit cost of a protocol $\Pi$ as $\mathsf{cost}_{\Pi}(G)=\lim_{k\to\infty}\tfrac{1}{k}\mathsf{cost}_{\Pi}(G,k)$ (this limit does not exist for all protocols, but it does in all the protocols we consider in the paper). As in the point-to-point communication model, a simple linear programming lower bound on the complexity of the problem can be obtained by noting that for every cut, the sum of the number of bits broadcast by the vertices incident to the cut has to be at least $k$. This implies that a natural lower bound on $\operatorname{\mathsf{OPT}}(G)$ is the maximum fractional packing (or minimum fractional transversal) of boundaries in the graph $G$, where a boundary is the set of vertices incident to the edges of a given cut in the graph.
@@
We note that it appears to be unknown whether $\operatorname{\mathsf{OPT}}(G)$ and $\operatorname{\mathsf{OPT}}(G,k)$ can be computed in polynomial time.
@@
We then design a new protocol in 2-connected graphs, which outperforms Theorem 1.2 on a number of natural graph classes. Our protocol uses ideas introduced by Alon, Efremenko and Sudakov [AES17] in the point-to-point model, relying on classical constructions in extremal combinatorics. Recall that a vertex cover in a graph $G$ is a subset of vertices intersecting every edge of $G$, and a total dominating set in $G$ is a subset $X$ of vertices such that every vertex in $G$ has a neighbor in $X$. The efficiency of our protocol depends on the minimum size of a total vertex cover in the graph $G$, denoted by $\mathsf{tvc}(G)$, where a total vertex cover is a subset of vertices which is both a vertex cover and a total dominating set. We prove the following result.
@@
Theorem 1.3 motivates the study of total vertex covers in graphs, for which we obtain new results which might be of independent interest. Our first application concerns the class of cycles. The linear programming lower bound on $\operatorname{\mathsf{OPT}}(C_{n})$ is of order $n/3$ (assigning weight $\tfrac{1}{3}$ to each vertex). On the other hand, as $\mathsf{wds}(C_{n})=\lfloor n/2\rfloor$, the simple protocol of Theorem 1.1 has per-bit cost $\lfloor n/2\rfloor$, and it can be checked that the protocol of Theorem 1.2 has per-bit cost $\tfrac{2}{3}n-\tfrac{2}{3}$ or $\tfrac{2}{3}n-\tfrac{1}{3}$, depending on the parity of $n$. As an immediate application of Theorem 1.3, we obtain an almost optimal protocol for cycles.
@@
It was proved in [CG17] that every cubic graph on $n$ vertices has a total vertex cover on at most $3n/4$ vertices. We extend this result to every degree by proving that for $d\geq 3$, every $d$-regular graph on $n$ vertices has a total vertex cover on at most $\tfrac{d}{d+1}\cdot n$ vertices. This is optimal for every $d$, since any total vertex cover of the complete graph $K_{d+1}$ (which is $d$-regular) contains at least $d$ vertices. It was proved in [CG17] that there are two more extremal connected graphs in the case $d=3$, and all the other connected cubic graphs have a total vertex cover with less than $3n/4$ vertices. Our result below implies that for $d\geq 5$, complete graphs are the only extremal examples.
@@ 2.1. Notation
In this paper, $\log$ denotes the binary logarithm. Let $G$ be a graph. For two disjoint sets $S,S^{\prime}\subseteq V(G)$, we denote by $E(S,S^{\prime})$ the set of edges of $G$ with the one endpoint in $S$ and the other one in $S^{\prime}$, and the cut defined by $S$ is $E(S,\bar{S})$, where $\bar{S}=V(G)\setminus S$. The boundary $B(S)$ of the cut defined by $S$ consists of all vertices incident to $E(S,\bar{S})$. Observe that for every $v\in V(G)$, the trivial boundary $B(\{v\})$ is precisely the closed neighborhood $N[v]$ of $v$ in $G$ (that is, $v$ together with its set of neighbors in $G$).
In the remainder we only consider connected graphs (otherwise the problem we consider has no solution, as vertices in different components cannot communicate to verify that their inputs are equal). A graph $G$ is 2-connected if it has at least 3 vertices and $G$ remains connected after the removal of any vertex.
2.2. A linear programming lower bound
@@ 4.1. The new protocol
Recall that a total vertex cover in a graph $G$ is a vertex subset that is both a total dominating set and a vertex cover. The minimum size of such a set is denoted by $\mathsf{tvc}(G)$. See Figures 2 and 3 for illustrations of total vertex covers in grids. In this section we prove Theorem 1.3, restated below as Theorem 4.1.
@@
In terms of transversals, a total vertex cover is an (integral) transversal of open neighborhoods (total dominating set) and of edges (vertex cover). As we are dealing with graphs with no isolated vertices, a total vertex cover always exists, for instance every vertex cover of such a graph $G$ is dominating and can be made total by adding a neighbor of each of its elements, so $\mathsf{tvc}(G)\leq 2\mathsf{vc}(G)$.
@@
See Figure 1 for an illustration (the three special copies of the triangle in the faithful host are depicted with different colors and width).
Figure 1. A faithful host of the triangle.
@@
Let $H$ be a 2-connected $n$-vertex graph and let $m$ be a positive integer. Then there is a faithful host for $H$ with classes of vertices $U_{1},\dots,U_{n}$, each of size $nm$, containing a least $m^{2}/e^{10\sqrt{\log m\log n}}$ special copies of $H$.
@@
Let $n$ denote the number of vertices of $G$, that we call $v_{1},\dots,v_{n}$, and let $m$ be the minimum integer such that $\frac{m^{2}}{e^{10\sqrt{\log m\log n}}}\geq 2^{k}$. By Lemma 4.3 there is a faithful host $F$ for $G$ with classes of vertices $U_{1},\dots,U_{n}$, each of size $nm$, containing a least $2^{k}$ special copies of $G$. For every $i\in\left\{1,\dots,n\right\}$, let us name $u_{i}^{1},\dots,u_{i}^{nm}$ the vertices of $U_{i}$. We associate to each $k$-bit word $w$ a special copy $G_{w}$ of $G$ in $F$ in an injective way (i.e. if $w\neq w^{\prime}$ then $G_{w}\neq G_{w^{\prime}}$). Since there are at least $2^{k}$ special copies, this is always possible.
@@
Protocol Communication: Each vertex $v$ of $S$ broadcasts its identity to its neighbors and the other vertices remain silent. Decision: A vertex $v$ of $G$ accepts the instance if and only if all identities received from its neighbors are consistent with its own identity.
@@ 4.2.3. Hypercubes
components, which concludes the proof of the claim. $\blacksquare$
@@ 4.2.4. Grids
In the $n\times n$ square grid $G=G_{n}$, consider the set $S$ of vertices $(i,j)\in[n]^{2}$ which either lie on the outerface, or are such that $i\in\{3j-2,3j-1\}\pmod{5}$ (in words, in row $j$ of the grid, we add to $S$ all vertices located in columns whose index is 3j23j-2 or 3j13j-1 modulo 5). Let $G_{S}$ be the subgraph of $G$ induced by the edges incident to $S$ (see Figure 2 for an illustration). Observe that $G_{S}$ is a spanning subgraph of $G$ and is 2-connected. By definition of $G_{S}$, it has a $S$ as a total vertex cover so $\mathsf{tvc}(G_{S})\leq|S|\leq\tfrac{2}{5}\,n^{2}+4n$. By Corollary 4.2, the protocol $\Pi_{2}$ has per-bit cost at most $\tfrac{1}{2}\,|S|\leq\tfrac{1}{5}\,n^{2}+2n$. By Lemma 2.2 we have $\operatorname{\mathsf{OPT}}(G)\geq\tfrac{1}{5}\,n^{2}$, and the result follows. ∎
@@ 4.2.5. Regular graphs
Every cubic graph on $n$ vertices has a total vertex cover on at most $3n/4$ vertices.
@@
By extending the proof of Theorem 4.9 we can prove that for $d\geq 3$, every $d$-regular graph on $n$ vertices has a total vertex cover on at most $\tfrac{d}{d+1}\cdot n$ vertices. This is optimal for every $d$, since any total vertex cover of the complete graph $K_{d+1}$ (which is $d$-regular) contains at least $d$ vertices. It was proved in [CG17] that there are two more extremal connected graphs in the case $d=3$, and all the other connected cubic graphs have a total vertex cover with less than $3n/4$ vertices. The result below implies that for $d\geq 5$, complete graphs are the only extremal examples.
@@ 5.2. Open problems
A natural problem is whether the fractional lower bound $\tau_{\rm{bnd}}^{*}(G)$ can always be attained by $\operatorname{\mathsf{OPT}}(G)$. We have not been able to find a graph for which the two parameters provably differ. A natural class to consider for potential counterexamples is the class of 2-edge-connected cubic graphs, which has a clean expression for $\tau_{\rm{bnd}}^{*}$. For this class we have only been able to find protocols of per-bit cost $3n/8$, while the fractional lower bound is $n/4$. Note that if we have $\operatorname{\mathsf{OPT}}(G)=\tau_{\rm{bnd}}^{*}(G)$ for every graph $G$, then $\operatorname{\mathsf{OPT}}(G)$ can indeed be computed in polynomial time.
@@
We have seen that for various types of grids we could obtain asymptotically optimal protocols for equality, using total vertex covers in some 2-connected subgraphs, of size at most twice the fractional lower bound. The three types of grids we have considered can all be seen as Cayley graphs of $\mathbb{Z}^{2}$, with different generating sets. As the three constructions are very similar, a natural question is whether our constructions can be generalized to any Cayley graph of $\mathbb{Z}^{2}$.
It remains an open question whether the upper bound of $4\operatorname{\mathsf{OPT}}(G)$ on the cost of the protocol $\Pi_{1}$ in Theorem 1.2 can be improved.
Finally, in this paper we have only considered protocols that are deterministic and static. It is an interesting problem to understand whether significantly better bounds can be obtained for protocols that are randomized and/or non-static.